# Line of Sight Bias in Dark Matter Inferences from Galaxy Cluster Mergers

### v1

@@ 4 Summary and discussion
We have found that current $\frac{\sigma_{\rm DM}}{m}$ constraints based on the scattering depth argument are on average biased because spherical symmetry is violated: the surface mass density along the LOS is lower than that experienced by DM particles as they pass along the merger axis. Because the surface mass density is inversely related to the inferred cross section, this yields an overestimate of the DM cross section—by a factor of ${\approx}2$ for surface mass densities measured along skewers (infinitesimally narrow paths through a halo). This bias is reduced to ${\approx}1.25$ for weak lensing mass measurements, which average sources over a large patch of sky. Note that this bias is distinct from the overall mass bias studied in Lee et al. (2023), which is smaller and stems from the overlap of two hypothetically spherical halos.
We studied dark matter only (DMO) halos; the presence of baryons will reduce bias by making halos rounder, but only slightly. For example, the simulations of Henson et al. (2016) found that the minor/major axis ratio shifts from $0.537$ for DMO to $0.576$ with baryons. This is only 40% of the scatter across halos. The intermediate/major axis ratio shifts by a similarly small amount when baryons are added.
@@
Although our argument was framed in terms of analytical approximations relating $\frac{\sigma_{\rm DM}}{m}$ to scattering depth (Markevitch et al., 2004; Harvey et al., 2014), the same bias applies to staged simulations based on the assumption of spherical symmetry. To overcome this bias, staged simulations should be staged with appropriate halo shapes and alignments, and we urge simulators to do so. A further step would be to go beyond the triaxial approximation and resimulate merging systems found in cosmological simulations. This would incorporate the effect of finer substructure and the fact that axis ratios vary with density (Jing & Suto, 2002). Resimulating appropriate boxes chosen from cosmological simulations would also naturally incorporate the appropriate range of halo alignments and pericenter distances.
We thank Rodrigo Stancioli for constructive suggestions. This work was supported by NSF grant number 2308383. The CosmoSim database used in this paper is a service by the Leibniz-Institute for Astrophysics Potsdam (AIP). The MultiDark database was developed in cooperation with the Spanish MultiDark Consolider Project CSD2009-00064. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) and the Partnership for Advanced Supercomputing in Europe (PRACE, www.prace-ri.eu) for funding the MultiDark simulation project by providing computing time on the GCS Supercomputer SuperMUC at Leibniz Supercomputing Centre (LRZ, www.lrz.de).

### v2

@@ 4 Summary and discussion
We have found that current $\frac{\sigma_{\rm DM}}{m}$ constraints based on the scattering depth argument are on average biased because spherical symmetry is violated: the surface mass density along the LOS is lower than that experienced by DM particles as they pass along the merger axis. Because the surface mass density is inversely related to the inferred cross section, this yields an overestimate of the DM cross section. We cast our results in terms of $\frac{\bar{\Sigma}_{\rm los}}{\bar{\Sigma}_{\rm CSV}}$, which in this argument is the same as the ratio of true to inferred DM cross section. We first introduced a thinking tool—skewers running through triaxial halo centers—that provides a strict lower bound to this ratio, independent of the density profile. In §3.2 we found that this ratio had a median value of 0.5 for typically used LOS, and with large scatter from halo to halo. Our weak lensing simulations (§2.4) show that in practice the ratio is about 0.8 for typically used LOS, and with less scatter. It also shows a modest dependence on the size of the region over which the surface mass density is averaged. Note that this bias is distinct from the overall mass bias studied in Lee et al. (2023), which is smaller and stems from the overlap of two hypothetically spherical halos.
We studied dark matter only (DMO) halos; the presence of baryons will reduce bias by making halos rounder, but only slightly. For example, the simulations of Henson et al. (2016) found that the median minor/major axis ratio shifts from $0.537$ for DMO to $0.576$ with baryons. This is only 40% of the scatter across halos. The intermediate/major axis ratio shifts by a similarly small amount when baryons are added. Still, there remains some uncertainty regarding the size of the effect given that different simulations with the same physics have different median axis ratios: the axis ratios for our BigMDPL halos are nontrivially lower than the DMO value in Henson et al. (2016). A comparable uncertainty is due to the size of the region within which $\bar{\Sigma}$ is measured (Figure 10), so further study is warranted—although some of these uncertainties will be obviated when DM constraints are derived from simulations as described in our final paragraph.
@@
Although our argument was framed in terms of analytical approximations relating $\frac{\sigma_{\rm DM}}{m}$ to scattering depth (Markevitch et al., 2004; Harvey et al., 2014), the same bias applies to staged simulations based on the assumption of spherical symmetry. To overcome this bias, staged simulations should be staged with appropriate halo shapes and alignments, and we urge simulators to do so. We checked the time evolution of the alignment, and we found that main halos are already well aligned at least ${\sim}1$ Gyr before pericenter. Hence one cannot rely on the alignment arising organically from processes (e.g. tidal effects) already present in the staged simulation. A further step would be to go beyond the triaxial approximation and resimulate merging systems found in cosmological simulations. This would incorporate the effect of finer substructure and the fact that axis ratios vary with density (Jing & Suto, 2002). Such simulations would also naturally incorporate the appropriate range of halo alignments and pericenter distances.
We thank the anonymous referee for numerous constructive suggestions. This work was supported by NSF grant number 2308383. The CosmoSim database used in this paper is a service by the Leibniz-Institute for Astrophysics Potsdam (AIP). The MultiDark database was developed in cooperation with the Spanish MultiDark Consolider Project CSD2009-00064. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) and the Partnership for Advanced Supercomputing in Europe (PRACE, www.prace-ri.eu) for funding the MultiDark simulation project by providing computing time on the GCS Supercomputer SuperMUC at Leibniz Supercomputing Centre (LRZ, www.lrz.de).